### v1

# The $u$-invariant of function fields in one variable

Abstract.
The $u$-invariant of a field is the largest dimension of an anisotropic quadratic torsion form over the field. In this article we obtain a bound on the $u$-invariant of function fields in one variable over a henselian valued field with arbitrary value group and with residue field of characteristic different from $2$. This generalises a theorem due to Harbater, Hartmann and Krashen and its extension due to Scheiderer. Their result covers the special case where the valuation is discrete.
Keywords: Quadratic form, henselian valued field, function field in one variable, quadratic form, local-global principle, field ordering, torsion form, general $u$-invariant
Classification (MSC 2020): 11E04, 11E81, 12D15, 12J10, 14H05
1. Introduction
The study of numerical field invariants in quadratic form theory was initiated by Irving Kaplansky in 1953. In [16], he introduces three quantities in ${}\mathbb{N}\cup\{\infty\}$ attached to an arbitrary field and suggests to study and compare them. One of them was later called the $u$-invariant, a term which goes back to Hanfried Lenz’ article [20] from 1963.
A field $K$ is called nonreal if $-1$ is a sum of squares in $K$, and otherwise $K$ is called real. For this introduction, let $K$ denote a nonreal field of characteristic different from $2$.
A quadratic form over $K$ is called universal if it represents all elements of $K$. Following Kaplansky and Lenz, the $u$-invariant of $K$ is defined as the smallest natural number $n$ such that any regular quadratic form in $n$ variables over $K$ is universal provided that such a natural number exists, and $\infty$ otherwise. This quantity in ${}\mathbb{N}\cup\{\infty\}$ is denoted by $u(K)$.
In the early 1970ies, Richard Elman and Tsit-Yuen Lam continued the study of this invariant in a series of articles. They modified the definition in a way that makes no difference for nonreal fields, but makes it more meaningful for real fields. We will recall their definition in Section 4.
As pointed out in [16, Theorem 3], one always has $u(K(\!(t)\!))=2u(K)$, where $K(\!(t)\!)$ denotes the field of Laurent series in $t$ over $K$. Consequently, and since $u(\mathbb{C})=1$, examples of fields of $u$-invariant $2^{r}$ for any given $r\in\mathbb{N}$ are easy to obtain. Incited by this observation, Kaplansky wonders in [16] whether $u(K)$ is always a power of $2$ or $\infty$. This question was answered in the negative in 1989 by Alexander Merkurjev. After giving in [24] a first construction of a field of $u$-invariant $6$, he showed in [25] that every positive even integer is the $u$-invariant of some field. Merkurjev’s construction of examples involves direct limits and produces fields which are not finitely generated over any proper subfield. When restricting to fields which are of finite transcendence degree over some familiar base field, no examples are known with a $u$-invariant that is not a power of $2$.
Let $n\in\mathbb{N}$ and let $F/K$ be a field extension of transcendence degree $n$. One may try to determine or at least bound the $u$-invariant of $F$ in terms of $n$ and properties of $K$. Well-known results of this sort are the following:
  • If $K$ is algebraically closed, then $u(F)\leqslant 2^{n}$, and equality holds if $F/K$ is finitely generated.
  • If $K$ is finite, then $u(F)\leqslant 2^{n+1}$, and equality holds if $F/K$ is finitely generated.
  • If $K$ is a complete discretely valued field with finite residue field, then $u(F)\leqslant 2^{n+2}$, and equality holds if $F/K$ is finitely generated.
In all three cases, the upper bounds are based on the study of systems of forms, while the lower bounds when $F/K$ is finitely generated follow from a classical argument using valuations. The first two results were established by Serge Lang in 1952 [18, Theorem 6, Cor.], whereas the last result was derived by David Leep in 2013 [19] from a result on systems of quadratic forms due to Roger Heath-Brown [14].
Note that in all the situations above, when $F/K$ is finitely generated, we have that $u(F)=2^{d}$ where $d$ is the cohomological dimension of $F$. This might suggest that $u(F)=2^{n+2}$ would hold as well when $K$ is a nonreal number field and $F/K$ a finitely generated extension of transcendence degree $n$. However, even for $n=1$ this is still an open problem. A more general open question is whether $u(F)$ can be bounded in terms of $u(K)$ and the transcendence degree of $F/K$.
We will now focus on function fields in one variable over a given base field. By a function field in one variable over $K$, we mean a finitely generated field extension of transcendence degree $1$. We can ask for sufficient conditions on $K$ for a certain number $u$ to serve as a common upper bound on $u(F)$ for all function fields in one variable $F/K$.
In particular, for any odd prime number $p$, applying 1.1 to the field of $p$-adic numbers $\mathbb{Q}_{p}$ with its natural valuation, one retrieves the result from [26] and [15] that $u(F)=8$ for every function field in one variable $F/\mathbb{Q}_{p}$.
Our main result provides a transfer principle for this common upper bound between a henselian valued field $K$ and its residue field, assuming the latter to be of characteristic different from $2$.
Our standard reference for the theory of valued fields is [13]. For a valuation $v$ on $K$, we denote by $Kv$ the residue field and by $vK$ the value group of $v$, and we call $v$ discrete if $vK\simeq\mathbb{Z}$. Given a valuation $v$ on $K$, we call the pair $(K,v)$ a valued field. A valuation $v$ on $K$ is called henselian if, for every finite field extension $L/K$, there is a unique extension of $v$ to a valuation $w$ on $L$; note that in this case $w$ is henselian, too. By [13, Theorem 1.3.1], every complete rank-$1$ valuation is henselian.
The characteristic of a ring $R$ is denoted by $\mathsf{char}(R)$. Our main result can be stated as follows:
Theorem 1.1 (6.2 & 6.9).
Assume that $(K,v)$ is a henselian valued field with $\mathsf{char}(Kv)\neq 2$. For $u\in\mathbb{N}$, the following are equivalent:
  1. $(i)$
    $u(F)\leqslant[vK:2vK]\cdot u$ for every function field in one variable $F/K$.
  2. $(ii)$
    $u(E)\leqslant u$ for every function field in one variable $E/Kv$.
Furthermore, if $u(E)=u$ for every function field in one variable $E/Kv$, then $u(F)=[vK:2vK]\cdot u$ for every function field in one variable $F/K$.
[…]
Unlike any previous results in the same direction, 1.1 holds for an arbitrary value group $vK$. Independently from our work, a weaker version of the implication from $(ii)$ to $(i)$ was obtained in [21, Theorem 4.1] for the case where $vK\subseteq\mathbb{R}$ (see also 6.3).
We further follow up on the step taken by Claus Scheiderer when extending [15, Cor. 4.12] to [28, Theorem 3] in making the statement more meaningful for real fields by using the definition of the $u$-invariant due to Elman and Lam, which we revisit in Section 4. 1.1 remains exactly the same also in that setting.
In the last part of Section 6, we also include some applications of our results.
An imminent challenge which we do not face in the present article is that of removing the hypothesis that the residue field has characteristic different from $2$. For the local-global principle from [23], this is perfectly possible, as is explained in [2, Theorem 7.1]. The study of quadratic forms over the completion $F^{w}$ of the function field $F$ with respect to a valuation $w$ on $F$, however, poses additional problems in residue characteristic $2$. We leave it to future study whether and how 1.1 can be extended to that case.
The study of quadratic forms over a complete valued field with residue characteristic different from $2$ and value group contained in $\mathbb{R}$ goes back to William Hetherington Durfee [9]. Some of the results were reproven for complete discrete valuations by Tonny Albert Springer [29]. The results from [9] extend to henselian valuations without any condition on the value group. While these statements and techniques are essentially well-known, to the best of our knowledge the literature does not cover them in the desirable generality or under minimal hypotheses. We therefore provide in Section 3 a self-contained and elementary presentation of those statements which are relevant to our topic. We indicate in 3.12 how this compares to the existing literature.
2. Valuations and torsion-free abelian groups
In the following statements we consider torsion-free abelian groups. We will denote the group operation additively.
Consider an abelian group $G$. The value $\mathsf{rrk}(G)=\mathsf{dim}_{\mathbb{Q}}(G\otimes_{\mathbb{Z}}\mathbb{Q})$ in ${}\mathbb{N}\cup\{\infty\}$ is called the rational rank of $G$. Hence, $\mathsf{rrk}(G)=|S|$ for any maximal $\mathbb{Z}$-linearly independent subset $S$ in $G$. In particular, $G$ is torsion if and only if $\mathsf{rrk}(G)=0$.
The following is an extension of [5, Lemma 3.4].
Proposition 2.1.
Let $G$ be a torsion-free abelian group and $H$ a subgroup of $G$ with $r=\mathsf{rrk}(G/H)<\infty$. Let $n\in\mathbb{N}^{+}$. Then $[G:nG]\leqslant n^{r}[H:nH]$. Furthermore, if $G/H$ is finitely generated, then $[G:nG]=n^{r}[H:nH]$.
Proof.
For any $s\in\mathbb{N}$ and elements $f_{1},\dots,f_{s}\in G$ representing different classes in $G/nG$, denoting by $F$ the subgroup of $G$ generated by $H\cup\{f_{1},\dots,f_{s}\}$, we have that $F/H$ is finitely generated and $f_{1},\dots,f_{s}$ represent distinct classes in $F/nF$, whereby $s\leqslant[F:nF]$. Therefore, to prove that $[G:nG]\leqslant n^{r}[H:nH]$, we may assume without loss of generality that $G/H$ is finitely generated.
We choose $g_{1},\dots,g_{r}\in G$ such that $(g_{i}+H\mid 1\leqslant i\leqslant r)$ is a maximal $\mathbb{Z}$-linearly independent family in $G/H$ and we denote by $G_{0}$ the subgroup of $G$ generated by $H\cup\{g_{1},\dots,g_{r}\}$. The choice of $g_{1},\dots,g_{r}$ implies that $G/G_{0}$ is torsion.
Assume that $G/H$ is finitely generated. Then $G/G_{0}$ is finite and $G_{0}\simeq\mathbb{Z}^{r}\times H$, whereby $[G_{0}:nG_{0}]=n^{r}[H:nH]$. Since $G$ is torsion-free, multiplication by $n$ induces an isomorphism $G/G_{0}\to nG/nG_{0}$. In particular $[G:G_{0}]=[nG:nG_{0}]$, whereby $[G_{0}:nG_{0}]=[G:nG]$. Therefore $[G:nG]=n^{r}[H:nH]$. ∎
By an ordered abelian group we mean an abelian group $(G,+,0)$ endowed with a total order relation $\leqslant$ such that, for any $x,y,z\in G$ with $x\leqslant y$, one has $x+z\leqslant y+z$. Note that any ordered abelian group is torsion-free as a group. When $(G,+,0,\leqslant)$ is an ordered abelian group, a convex subgroup is a subgroup $H$ of $G$ such that for all $x\in H$ and $y\in G$ with $0\leqslant y\leqslant x$, we have $y\in H$. One sees that the set of convex subgroups of $G$ is totally ordered by inclusion. The cardinality of the set of proper convex subgroups of $G$ will be called the rank of $G$. By [13, Prop. 3.4.1], the rank of $G$ is bounded from above by $\mathsf{rrk}(G)$.
Given a valuation $v$ on a field $K$, we refer to the rank of the value group $vK$ also as the rank of $v$. Note that an ordered abelian group can be embedded into $\mathbb{R}$ if and only if it is of rank at most $1$. (Consequently, the rank-$1$ valuations on $K$ correspond precisely to the non-trivial absolute values on $K$.)
When $L/K$ is an extension of fields and $v$ and $w$ are valuations on $K$ and $L$ respectively, we say that $w$ extends $v$ if $vK\subseteq wL$ and $v=w|_{K}$, and we call $(L,w)/(K,v)$ an extension of valued fields.
Proposition 2.2.
Let $F/K$ be a function field in one variable and let $w$ be a valuation on $F$ and $v=w|_{K}$. Then one of the following holds:
  1. $(1)$
    $Fw/Kv$ is a finite field extension and $wF/vK$ is a finitely generated $\mathbb{Z}$-module with $\mathsf{rrk}(wF/vK)=1$.
  2. $(2)$
    $Fw/Kv$ is a function field in one variable and $wF/vK$ is finite.
  3. $(3)$
    $Fw/Kv$ is an algebraic field extension and $wF/vK$ is a torsion group.
Proof.
This follows from [13, Theorem 3.4.3]. ∎
Lemma 2.3.
Let $(K,v)$ be a henselian valued field and $K_{0}$ a subfield of $K$. There exists a subfield $K_{1}$ of $K$ containing $K_{0}$ such that $vK_{1}/vK_{0}$ is torsion, $vK/vK_{1}$ is torsion-free, $Kv/K_{1}v$ is algebraic and purely inseparable and $v|_{K_{1}}$ is henselian.
Proof.
Using Zorn’s Lemma, we choose $K_{1}$ as an extension of $K_{0}$ inside $K$ which is maximal with respect to the property that the quotient $vK_{1}/vK_{0}$ is torsion.
Consider $t\in K$ algebraic over $K_{1}$. Then ${[vK_{1}(t):vK_{1}]\leqslant[K_{1}(t):K_{1}]<\infty}$, so $vK_{1}(t)/vK_{1}$ is torsion, and it follows by the choice of $K_{1}$ that $t\in K_{1}$. Therefore, $K_{1}$ is relatively algebraically closed in $K$.
Consider $t\in K$ transcendental over $K_{1}$. Then $K_{1}(t)/K_{1}$ is a rational function field in one variable, and hence it follows by 2.2 from the choice of $K_{1}$ that the extension $K_{1}(t)v/K_{1}v$ is finite and $vK_{1}(t)/vK_{1}\simeq\mathbb{Z}$. Having this for any $t\in K$, we conclude that the extension $Kv/K_{1}v$ is algebraic and $vK/vK_{1}$ is torsion-free.
Since $(K,v)$ is henselian and $K_{1}$ is relatively algebraically closed in $K$, it follows that $(K_{1},v|_{K_{1}})$ is henselian as well.
Consider $\xi\in Kv$ separable over $K_{1}v$. Since $(K,v)$ is henselian, $\xi$ is the residue of some element $t\in\mathcal{O}_{v}$ algebraic over $K_{1}$. Since $K_{1}$ is algebraically closed in $K$, we obtain that $t\in K_{1}$, whereby $\xi\in K_{1}v$. This shows that $Kv/K_{1}v$ is purely inseparable. ∎
3. Valued fields and residue forms
We assume familiarity with standard facts from quadratic form theory over fields as covered in particular by the first chapters of [17] and [10]. On this basis, we will give a succinct introduction to the theory of quadratic forms over valued fields with residue characteristic different from $2$. While in the case of a complete discrete valuation, all these results are well-known, we have not found a presentation of this theory in the literature which would provide the statements in this generality. Our presentation resembles that of [9], a pioneer article on the topic of quadratic forms over valuation rings. The terminology used in [9] is meanwhile a bit off the standard, and the statements in [9] are formulated for a valuation that is assumed to be complete of rank-$1$, while the arguments mostly work in more generality. Our presentation highlights that large parts of the theory may be developed without recourse to a henselization or completion of the valuation ring. To introduce the concepts and formulate a first few facts, we deal with quadratic forms over a local ring in which $2$ is invertible.
Let $R$ be a commutative ring. A quadratic form over $R$ is a homogeneous polynomial $\varphi$ of degree $2$ in a prescribed finite tuple of variables whose length is called the dimension of $\varphi$ and denoted by $\mathsf{dim}(\varphi)$. A quadratic form $\varphi$ over $R$ gives rise to a map $R^{n}\to R$ for $n=\mathsf{dim}(\varphi)$, defined by evaluating $\varphi$ on $n$-tuples, as well as to an $R$-bilinear form
$\mathfrak{b}_{\varphi}:R^{n}\times R^{n}\to R:(x,y)\mapsto\varphi(x+y)-\varphi(x)-\varphi(y).$
Let $\varphi$ be a quadratic form over $R$. Let $n=\mathsf{dim}(\varphi)$ and $X=(X_{1},\dots,X_{n})$. We say that $\varphi$ represents $a\in R$ if $a=\varphi(x)$ for some $x\in R^{n}$. We call $\varphi$ isotropic if $\varphi(x)=0$ for some $x=(x_{1},\dots,x_{n})\in R^{n}$ such that $R$ is generated as an ideal by the components of $x$. (If $R$ is a local ring, then the latter means that $x_{i}\in{R}^{\times}$ for some $i\in\{1,\dots,n\}$.) If $\varphi$ is not isotropic, then we call $\varphi$ anisotropic.
Let $n\in\mathbb{N}$ and $X=(X_{1},\dots,X_{n})$. Given elements $a_{1},\dots,a_{n}\in R$, we denote the $n$-dimensional quadratic form $a_{1}X_{1}^{2}+\ldots+a_{n}X_{n}^{2}$ over $R$ by $\langle a_{1},\dots,a_{n}\rangle_{R}$, or simply by $\langle a_{1},\dots,a_{n}\rangle$ if the ring $R$ is clear from the context. A quadratic form of this shape (that is, where the terms $X_{i}X_{j}$ with $i\neq j$ all have coefficient 00) is called diagonalized or a diagonal form. Let $\varphi,\psi\in R[X]$ be $n$-dimensional quadratic forms over $R$. We call $\varphi$ and $\psi$ isometric if $\psi(X)=\varphi(X\cdot\alpha)$ for some invertible matrix $\alpha\in\mathbb{M}_{n}(R)$, and we write $\varphi\simeq\psi$ to indicate this. If $\varphi$ and $\psi$ are quadratic forms over $R$, $m=\mathsf{dim}(\varphi)$ and $n=\mathsf{dim}(\psi)$, then we denote by $\varphi\perp\psi$ the $(m+n)$-dimensional quadratic form $\varphi(X_{1},\dots,X_{m})+\psi(X_{m+1},\dots,X_{m+n})$ over $R$. We further denote by $\mathbb{H}_{R}$ or by $\mathbb{H}$ the quadratic form $X_{1}X_{2}$ over $R$.
The next few observations are under the assumption that $2\in{R}^{\times}$. This implies that $\mathbb{H}\simeq_{R}\langle 1,-1\rangle$. Consider an $n$-dimensional quadratic form $\varphi$ in $X=(X_{1},\dots,X_{n})$ over $R$. Then $\varphi=XAX^{\mathsf{t}}$ for a unique symmetric matrix $A\in\mathbb{M}_{n}(R)$. We set $\mathsf{det}(\varphi)=\mathsf{det}(A)$ and call this the determinant of $\varphi$. We call $\varphi$ regular if $A$ is invertible in $\mathbb{M}_{n}(R)$, which is if and only if $\mathsf{det}(A)\in{R}^{\times}$. Note that $\mathsf{det}(\varphi\perp\psi)=\mathsf{det}(\varphi)\cdot\mathsf{det}(\psi)$ for any quadratic form $\psi$ over $R$. Clearly, if $\varphi$ is regular, then so is every quadratic form isometric to $\varphi$.
While our main attention is to the case where $R$ is a valuation ring with $2\in{R}^{\times}$, the next few statements hold more generally when $R$ is local and $2\in{R}^{\times}$.
Proposition 3.1.
Assume that $R$ is local and $2\in{R}^{\times}$. Let $\varphi$ be a quadratic form over $R$ and $b\in{R}^{\times}$. Then $\varphi$ represents $b$ if and only if $\varphi\simeq\langle b\rangle_{R}\perp\psi$ for a quadratic form $\psi$ over $R$.
Proof.
Clearly, $b$ is represented by $\langle b\rangle_{R}$ and hence also by $\varphi$ if $\varphi\simeq\langle b\rangle_{R}\perp\psi$ for some quadratic form $\psi$ over $R$.
To prove the converse, assume that $\varphi(x)=b$ for some $x=(x_{1},\dots,x_{n})\in R^{n}$. Since $b\in R^{\times}$ and $R$ is local, there exists $i\in\{1,\ldots,n\}$ with $x_{i}\in R^{\times}$. We may without loss of generality assume that $x_{1}\in R^{\times}$.
Let $(e_{1},\ldots,e_{n})$ be the canonical basis of $R^{n}$. We set $e_{1}^{\prime}=x$ and further $e_{i}^{\prime}=e_{i}-(2b)^{-1}\mathfrak{b}_{\varphi}(e_{1}^{\prime},e_{i})e_{1}^{\prime}\in R^{n}$ for $2\leqslant i\leqslant n$. Then $(e_{1}^{\prime},\dots,e_{n}^{\prime})$ is also an $R$-basis of $R^{n}$. Let $\alpha\in\mathbb{M}_{n}(R)$ be the matrix of the base change from $(e_{1}^{\prime},\ldots,e_{n}^{\prime})$ to $(e_{1},\ldots,e_{n})$. Then $\alpha$ is invertible in $\mathbb{M}_{n}(R)$. Using that $\mathfrak{b}_{\varphi}$ is $R$-bilinear, one computes that $\mathfrak{b}_{\varphi}(e_{1}^{\prime},e_{i}^{\prime})=0$ for $2\leqslant i\leqslant n$. Thus, letting $X=(X_{1},\dots,X_{n})$, we have that $\varphi(X)\simeq\varphi(X\alpha)=bX_{1}^{2}+\psi(X_{2},\dots,X_{n})$ for an $(n-1)$-dimensional quadratic form $\psi$ over $R$. ∎
Corollary 3.2.
Assume that $R$ is local and $2\in{R}^{\times}$. Let $\varphi$ be a quadratic form over $R$ and $n=\mathsf{dim}(\varphi)$. Then $\varphi$ is regular if and only if $\varphi\simeq\langle a_{1},\dots,a_{n}\rangle_{R}$ for certain $a_{1},\dots,a_{n}\in{R}^{\times}$.
Proof.
Clearly, if $\varphi\simeq\langle a_{1},\dots,a_{n}\rangle_{R}$ for certain $a_{1},\dots,a_{n}\in{R}^{\times}$, then $\varphi$ is regular.
The converse is proven by induction on $n$. For $n=0$, the statement holds trivially. Assume that $n>0$. Let $\mathfrak{m}$ be the maximal ideal of $R$. Then ${R}^{\times}=R\smallsetminus\mathfrak{m}$ and, since $\varphi$ is regular, the entries of the associated matrix cannot all lie in $\mathfrak{m}$. So there exist $x,y\in R^{n}$ with $\mathfrak{b}_{q}(x,y)\notin\mathfrak{m}$. As $\mathfrak{b}_{q}(x,y)=\varphi(x+y)-\varphi(x)-\varphi(y)$, we conclude that $\varphi(z)\notin\mathfrak{m}$ for some $z\in\{x,y,x+y\}$. Letting $a_{1}=\varphi(z)$, we have that $a_{1}\in R\smallsetminus\mathfrak{m}={R}^{\times}$ and obtain by 3.1 that $\varphi\simeq\langle a_{1}\rangle\perp\psi$ for some $(n-1)$-dimensional quadratic form $\psi$ over $R$. As $\varphi$ is regular, so is $\langle a_{1}\rangle\perp\psi$. Hence $a_{1}\mathsf{det}(\psi)\in{R}^{\times}$, which implies that $\mathsf{det}(\psi)\in{R}^{\times}$. Hence $\psi$ is regular. By the induction hypothesis, we obtain that $\psi\simeq\langle a_{2},\dots,a_{n}\rangle_{R}$ for certain $a_{2},\dots,a_{n}\in{R}^{\times}$, whereby $\varphi\simeq\langle a_{1},\dots,a_{n}\rangle_{R}$. ∎
Lemma 3.3.
Assume that $R$ is local with $2\in{R}^{\times}$. Let $\mathfrak{m}$ denote the maximal ideal of $R$. Let $x\in{R}^{\times}$ and $y,c\in R$. If $|R/\mathfrak{m}|=3$, then assume that $c\notin 1+\mathfrak{m}$. Then there exist $x^{\prime}\in R$ and $y^{\prime}\in{R}^{\times}$ such that $x^{2}-cy^{2}=x^{\prime 2}-cy^{\prime 2}$.
Proof.
If $y\in{R}^{\times}$, then we may take $x^{\prime}=x$ and $y^{\prime}=y$. Assume now that $y\notin{R}^{\times}$. In view of the hypotheses, there exists $\lambda\in R$ such that $1-\lambda^{2}c\in{R}^{\times}$ and $2\lambda x+(1+\lambda^{2}c)y\in{R}^{\times}$. Set $x^{\prime}=(1-\lambda^{2}c)^{-1}((1+\lambda^{2}c)x+2\lambda cy)$ and $y^{\prime}=(1-\lambda^{2}c)^{-1}(2\lambda x+(1+\lambda^{2}c)y)$. Then
$(1-\lambda^{2}c)^{2}(x^{\prime 2}-cy^{\prime 2})$ $=$ $((1+\lambda^{2}c)x+2\lambda cy)^{2}-c(2\lambda x+(1+\lambda^{2}c)y)^{2}$
$=$ $(1-\lambda^{2}c)^{2}(x^{2}-cy^{2})\,,$
whereby $x^{2}-cy^{2}=x^{\prime 2}-cy^{\prime 2}$. ∎
For $n\in\mathbb{N}$, a vector $x\in R^{n}$ is called primitive if its coordinates generate $R$ as an ideal; as $R$ is local, this just says that one of the coordinates of $x$ lies in ${R}^{\times}$.
For valuation rings, the following statement is [6, Theorem 4.5], where it is attributed to J.-L. Colliot-Thélène and M. Kneser. We give a different proof.
Proposition 3.4.
Assume that $R$ is local with $2\in{R}^{\times}$. Let $\varphi$ be a regular quadratic form over $R$, $n=\mathsf{dim}(\varphi)$ and $b\in R$. If $x_{0}^{2}b=\varphi(x)$ for some $x_{0}\in R$ and some primitive vector $x\in R^{n}$, then $b$ is represented by $\varphi$.
Proof.
Let $\mathfrak{m}$ denote the maximal ideal of $R$. By 3.2, we may assume without loss of generality that $\varphi=\langle a_{1},\dots,a_{n}\rangle_{R}$ with $a_{1},\dots,a_{n}\in{R}^{\times}$. Let $x=(x_{1},\dots,x_{n})\in R^{n}$ be primitive and $x_{0}\in R$ such that $x_{0}^{2}b=\varphi(x)$. We may permute $a_{1},\dots,a_{n}$ and assume without loss of generality that, for some $m\in\{1,\dots,n\}$, we have $x_{1},\dots,x_{m}\in{R}^{\times}$ and $x_{m+1},\dots,x_{n}\in\mathfrak{m}$.
Assume first that $|R/\mathfrak{m}|>3$ or $a_{1}\not\equiv b\,\,\mathsf{mod}\,\,\mathfrak{m}$. Set $c=a_{1}^{-1}b$. Using 3.3, we obtain that $x_{1}^{2}-cx_{0}^{2}=z_{1}^{2}-cz_{0}^{2}$ for certain $z_{0}\in{R}^{\times}$ and $z_{1}\in R$. We set $y_{1}=z_{0}^{-1}z_{1}$ to obtain that
$a_{2}x_{2}^{2}+\dots+a_{n}x_{n}^{2}=bx_{0}^{2}-a_{1}x_{1}^{2}=bz_{0}^{2}-a_{1}z_{1}^{2}=z_{0}^{2}(b-a_{1}y_{1}^{2})\,.$
Letting $y_{i}=z_{0}^{-1}x_{i}$ for $2\leqslant i\leqslant n$, we obtain that $b=a_{1}y_{1}^{2}+\ldots+a_{n}y_{n}^{2}$, as desired.
Consider now the case where $|R/\mathfrak{m}|=3$ and $a_{1}\equiv b\,\,\mathsf{mod}\,\,\mathfrak{m}$. If $x_{0}\in{R}^{\times}$, then we may set $y_{i}=x_{0}^{-1}x_{i}$ for $1\leqslant i\leqslant n$ to obtain the desired equality. Assume now that $x_{0}\in\mathfrak{m}$. Since $a_{1}x_{1}^{2}\in{R}^{\times}$ and $a_{1}x_{1}^{2}+\ldots+a_{n}x_{n}^{2}=bx_{0}^{2}\in\mathfrak{m}$, we obtain that $m\geqslant 2$, whereby $x_{2}\in{R}^{\times}$. If $a_{2}\not\equiv b\,\,\mathsf{mod}\,\,\mathfrak{m}$, then we switch the roles of $a_{1}$ and $a_{2}$ and conclude by the previous case. So we are left to consider the case where $a_{1}\equiv a_{2}\equiv b\,\,\mathsf{mod}\,\,\mathfrak{m}$. We set $c=a_{1}x_{1}^{2}+a_{2}x_{2}^{2}$. Then $c\equiv-b\not\equiv b\,\,\mathsf{mod}\,\,\mathfrak{m}$, and we conclude by the previous case that there exist $z,y_{3},y_{4},\ldots,y_{n}\in R$ such that $b=cz^{2}+a_{3}y_{3}^{2}+\ldots+a_{n}y_{n}^{2}$. Then letting $y_{1}=zx_{1}$ and $y_{2}=zx_{2}$, we obtain that $b=a_{1}y_{1}^{2}+\ldots+a_{n}y_{n}^{2}$. ∎
In the sequel, let $K$ be a field and $v$ a valuation on $K$. Recall that we denote by $\mathcal{O}_{v}$ the valuation ring of $v$, by $\mathfrak{m}_{v}$ its maximal ideal, by $Kv$ the residue field $\mathcal{O}_{v}/\mathfrak{m}_{v}$ and by $vK$ the value group $v({K}^{\times})$. We assume in the sequel that $\mathsf{char}(Kv)\neq 2$. This means that $2\in{\mathcal{O}}^{\times}_{v}$. In particular, $\mathsf{char}(K)\neq 2$.
Proposition 3.5.
Let $n\in\mathbb{N}$ and $a_{1},\dots,a_{n},b\in{\mathcal{O}}^{\times}_{v}$.
  1. $(a)$
    If $\langle a_{1},\dots,a_{n}\rangle_{K}$ represents $b$, then $n\geqslant 1$ and there exist $b_{2},\dots,b_{n}\in{\mathcal{O}}^{\times}_{v}$ such that $\langle a_{1},\dots,a_{n}\rangle_{\mathcal{O}_{v}}\simeq\langle b,b_{2},\dots,b_{n}\rangle_{\mathcal{O}_{v}}$.
  2. $(b)$
    If $\langle a_{1},\dots,a_{n}\rangle_{K}$ is isotropic, then $n\geqslant 2$ and there exist $b_{3},\dots,b_{n}\in{\mathcal{O}}^{\times}_{v}$ such that $\langle a_{1},\dots,a_{n}\rangle_{\mathcal{O}_{v}}\simeq\mathbb{H}_{\mathcal{O}_{v}}\perp\langle b_{3},\dots,b_{n}\rangle_{\mathcal{O}_{v}}$.
Proof.
We set $\varphi=\langle a_{1},\dots,a_{n}\rangle_{\mathcal{O}_{v}}$ and $\varphi_{K}=\langle a_{1},\dots,a_{n}\rangle_{K}$.
$(a)$ Assume that $\varphi_{K}$ represents $b$. Then $bx_{0}^{2}=a_{1}x_{1}^{2}+\ldots+a_{n}x_{n}^{2}$ for certain $x_{0},\dots,x_{n}\in\mathcal{O}_{v}$, not all contained in $\mathfrak{m}_{v}$. It follows that $x_{i}\in{\mathcal{O}}^{\times}_{v}$ for some $i\in\{1,\dots,n\}$. We obtain by 3.4 that $b=a_{1}y_{1}^{2}+\ldots+a_{n}y_{n}^{2}$ for certain $y_{1},\dots,y_{n}\in\mathcal{O}_{v}$. We conclude by 3.1 that $\varphi\simeq\langle b\rangle_{\mathcal{O}_{v}}\perp\psi$ for some $(n-1)$-dimensional quadratic form $\psi$ over $\mathcal{O}_{v}$. Using 3.2 we obtain that $\psi\simeq\langle b_{2},\dots,b_{n}\rangle_{\mathcal{O}_{v}}$ for certain $b_{2},\dots,b_{n}\in{\mathcal{O}}^{\times}_{v}$. Then $\varphi\simeq\langle b,b_{2},\dots,b_{n}\rangle_{\mathcal{O}_{v}}$.
$(b)$ Assume that $\varphi_{K}$ is isotropic. Hence $a_{1}x_{1}^{2}+\ldots+a_{n}x_{n}^{2}=0$ for certain $x_{1},\dots,x_{n}\in K$ which are not all zero. If $x_{1}=0$, then $\langle a_{2},\dots,a_{n}\rangle_{K}$ is isotropic. As every isotropic form over $K$ represents all elements of $K$, we obtain in any case that $-a_{1}$ is represented by $\langle a_{2},\dots,a_{n}\rangle_{K}$. Applying part $(a)$ to $\langle a_{2},\dots,a_{n}\rangle_{K}$, we obtain that $\langle a_{2},\dots,a_{n}\rangle_{\mathcal{O}_{v}}\simeq\langle-a_{1},b_{3},\dots,b_{n}\rangle_{\mathcal{O}_{v}}$ for certain $b_{3},\dots,b_{n}\in{\mathcal{O}}^{\times}_{v}$. As $\langle a_{1},-a_{1}\rangle_{\mathcal{O}_{v}}\simeq\mathbb{H}_{\mathcal{O}_{v}}$, we conclude that $\varphi\simeq\mathbb{H}_{\mathcal{O}_{v}}\perp\langle b_{3},\dots,b_{n}\rangle_{\mathcal{O}_{v}}$. ∎
A quadratic form $\varphi$ over $K$ is called $v$-unimodular if $\varphi\simeq\langle a_{1},\dots,a_{n}\rangle_{K}$ for $n=\mathsf{dim}(\varphi)$ and certain $a_{1},\dots,a_{n}\in{\mathcal{O}}^{\times}_{v}$.
Corollary 3.6.
A quadratic form over $K$ is $v$-unimodular if and only if its anisotropic part is $v$-unimodular.
Proof.
This follows by induction from 3.5 $(b)$. ∎
Proposition 3.7.
Let $n\in\mathbb{N}$. Let $a_{1},\dots,a_{n},c_{1},\dots,c_{n}\in{\mathcal{O}}^{\times}_{v}$. Assume that we have $\langle a_{1},\dots,a_{n}\rangle_{K}\simeq\langle c_{1},\dots,c_{n}\rangle_{K}$. Then $\langle a_{1},\dots,a_{n}\rangle_{\mathcal{O}_{v}}\simeq\langle c_{1},\dots,c_{n}\rangle_{\mathcal{O}_{v}}$ and $\langle\overline{a}_{1},\dots,\overline{a}_{n}\rangle_{Kv}\simeq\langle\overline{c}_{1},\dots,\overline{c}_{n}\rangle_{Kv}$.
Proof.
An isometry over $\mathcal{O}_{v}$ induces an isometry over $Kv$, so we only need to prove the first part. The proof is by induction on $n$. For $n=0$ there is nothing to prove. Assume that $n\geqslant 1$. As $c_{1}$ is represented by $\langle a_{1},\dots,a_{n}\rangle_{K}$, 3.5 yields that $\langle a_{1},\dots,a_{n}\rangle_{\mathcal{O}_{v}}\simeq\langle c_{1},b_{2},\dots,b_{n}\rangle_{\mathcal{O}_{v}}$ for certain $b_{2},\dots,b_{n}\in{\mathcal{O}}^{\times}_{v}$. Then $\langle c_{1},\dots,c_{n}\rangle_{K}\simeq\langle a_{1},\dots,a_{n}\rangle_{K}\simeq\langle c_{1},b_{2},\dots,b_{n}\rangle_{K}$. Witt’s Cancellation Theorem [17, Chap. I, Theorem 4.2] yields that $\langle c_{2},\dots,c_{n}\rangle_{K}\simeq\langle b_{2},\dots,b_{n}\rangle_{K}$. Now, we obtain from the induction hypothesis, that $\langle c_{2},\dots,c_{n}\rangle_{\mathcal{O}_{v}}\simeq\langle b_{2},\dots,b_{n}\rangle_{\mathcal{O}_{v}}$. Then $\langle a_{1},\dots,a_{n}\rangle_{\mathcal{O}_{v}}\simeq\langle c_{1},b_{2},\dots,b_{n}\rangle_{\mathcal{O}_{v}}\simeq\langle c_{1},c_{2}\dots,c_{n}\rangle_{\mathcal{O}_{v}}$. ∎
Assume that $\varphi$ is $v$-unimodular. Let $n=\mathsf{dim}(\varphi)$ and $a_{1},\dots,a_{n}\in{\mathcal{O}}^{\times}_{v}$ such that $\varphi\simeq\langle a_{1},\dots,a_{n}\rangle_{K}$. We associate to $\varphi$ the quadratic form $\overline{\varphi}^{v}=\langle\overline{a}_{1},\dots,\overline{a}_{n}\rangle_{Kv}$ over $Kv$. By 3.7, the quadratic form $\overline{\varphi}^{v}$ is uniquely determined up to isometry by $\varphi$. We call $\overline{\varphi}^{v}$ the residue form of $\varphi$ (with respect to $v$). We often simply write $\overline{\varphi}$ for $\overline{\varphi}^{v}$, if the context clarifies the reference to the valuation $v$.
Corollary 3.8.
Let $\varphi$ be a $v$-unimodular quadratic form over $K$. If $\varphi$ is isotropic, then $\overline{\varphi}^{v}$ is isotropic.
Proof.
This follows directly from 3.5 $(b)$. ∎
Proposition 3.9.
Let $\varphi$ be a quadratic form over $K$. Let $\mathcal{C}$ be a set of representatives of ${K}^{\times}/v^{-1}(2vK)$. There exist $r\in\mathbb{N}$, distinct $c_{1},\dots,c_{r}\in\mathcal{C}$, and $v$-unimodular quadratic forms $\varphi_{1},\dots,\varphi_{r}$ over $K$ such that $\varphi\simeq c_{1}\varphi_{1}\perp\ldots\perp c_{r}\varphi_{r}$.
Proof.
By the hypothesis on $\mathcal{C}$ we have $vK=\bigcup_{c\in\mathcal{C}}(v(c)+2vK)$. We fix a diagonalization of $\varphi$ and then choose $c_{1},\dots,c_{r}\in\mathcal{C}$ such that $\bigcup_{i=1}^{r}(v(c_{i})+2vK)$ contains the values of all entries of this diagonalization. For $i\in\{1,\dots,r\}$ and $a\in{K}^{\times}$ with $v(a)\in v(c_{i})+2vK$, we have $\langle a\rangle\simeq\langle c_{i}u\rangle=c_{i}\langle u\rangle$ for some $u\in{\mathcal{O}}^{\times}_{v}$. Hence, any of the diagonal entries can be replaced by $c_{i}u$ for some $i\in\{1,\dots,r\}$ and $u\in{\mathcal{O}}^{\times}_{v}$. This leads to a presentation $\varphi\simeq c_{1}\varphi_{1}\perp\ldots\perp c_{r}\varphi_{r}$ where $\varphi_{1},\dots,\varphi_{r}$ are $v$-unimodular quadratic forms over $K$. ∎
We come to the main statements of this section. They show that, if the valuation $v$ on $K$ is henselian, then unimodular quadratic forms over $K$ are characterised by their residue forms. We will formulate those statements under the hypothesis that $1+\mathfrak{m}_{v}\subseteq{K}^{\times 2}$, which is in particular implied when $v$ is henselian (using that $\mathsf{char}(Kv)\neq 2$); see e.g. [13, Cor. 4.2.4].
Theorem 3.10.
Assume that $1+\mathfrak{m}_{v}\subseteq{K}^{\times 2}\!$. Let $\varphi$ be a $v$-unimodular quadratic form over $K$ and $b\in{\mathcal{O}}^{\times}_{v}$. The following equivalences hold:
  1. $(a)$
    $\varphi$ represents $b$ if and only if $\overline{\varphi}^{v}$ represents $\overline{b}$.
  2. $(b)$
    $\varphi$ is isotropic if and only if $\overline{\varphi}^{v}$ is isotropic.
  3. $(c)$
    $\varphi$ is hyperbolic if and only if $\overline{\varphi}^{v}$ is hyperbolic.
Proof.
Let $n\in\mathbb{N}$ and $a_{1},\dots,a_{n}\in{\mathcal{O}}^{\times}_{v}$ be such that $\varphi\simeq\langle a_{1},\dots,a_{n}\rangle_{K}$, whereby $\overline{\varphi}\simeq\langle\overline{a}_{1},\dots,\overline{a}_{n}\rangle_{Kv}$.
$(a)$ If $\varphi$ represents $b$ over $K$, then it represents $b$ over $\mathcal{O}_{v}$, by 3.5, and hence $\overline{\varphi}^{v}$ represents $\overline{b}$ over $Kv$. For the converse, assume that $\overline{\varphi}^{v}$ represents $\overline{b}$ over $Kv$. Hence, there exist $x\in\mathcal{O}_{v}^{n}$ such that $\varphi(x)\equiv b\,\,\mathsf{mod}\,\,\mathfrak{m}_{v}$. It follows that $\varphi(x)b^{-1}\in 1+\mathfrak{m}_{v}\subseteq{K}^{\times 2}$, whereby $\varphi(x){K}^{\times 2}=b{K}^{\times 2}$. Hence $\varphi$ represents $b$.
$(b)$ If $\varphi$ is isotropic, then so is $\overline{\varphi}^{v}$, by 3.8. If $\overline{\varphi}$ is isotropic, then $\overline{\varphi}^{\prime}=\langle\overline{a}_{2},\dots,\overline{a}_{n}\rangle_{Kv}$ represents $-\overline{a}_{1}$, and it follows by $(a)$ that $\varphi^{\prime}$ represents $-a_{1}$, whereby $\varphi$ is isotropic.
$(c)$ We prove the equivalence by induction on $n$. If $n=0$, then $\varphi$ and $\overline{\varphi}^{v}$ are trivially hyperbolic. Assume that $n>0$. If $\varphi$ is anisotropic, then so is $\overline{\varphi}^{\prime}$, by $(b)$, and as $n>0$, it follows that neither $\varphi$ nor $\overline{\varphi}$ is hyperbolic. Assume that $\varphi$ is isotropic. By 3.5 $(b)$, there exist $b_{3},\dots,b_{n}\in{\mathcal{O}}^{\times}_{v}$ such that $\varphi\simeq\mathbb{H}_{K}\perp\langle b_{3},\dots,b_{n}\rangle_{K}$ and $\overline{\varphi}\simeq\mathbb{H}_{Kv}\perp\langle\overline{b}_{3},\dots,\overline{b}_{n}\rangle_{Kv}$. Now, by Witt’s Cancellation Theorem [17, Chap. I, Theorem 4.2], $\varphi$ is hyperbolic if and only if $\rho=\langle b_{3},\dots,b_{n}\rangle_{K}$ is hyperbolic, and $\overline{\varphi}^{v}$ is hyperbolic if and only if $\overline{\rho}^{v}=\langle\overline{b}_{3},\dots,\overline{b}_{n}\rangle_{Kv}$ is hyperbolic. By the induction hypothesis, $\rho$ is hyperbolic if and only if $\overline{\rho}^{v}$ is hyperbolic. From this together we conclude that $\varphi$ is hyperbolic if and only if $\overline{\varphi}^{v}$ is hyperbolic. ∎
Let $\varphi$ be a quadratic form over $K$. For $m\in\mathbb{N}$, we denote by $m\times\varphi$ the $m$-fold orthogonal sum $\varphi\perp\ldots\perp\varphi$. We say that $\varphi$ is torsion if $m\times\varphi$ is hyperbolic for some $m\in\mathbb{N}^{+}$.
Part of the following statement recovers [9, Theorem 1] and its corollary.
Theorem 3.11 (Durfee).
[…]
Proof.
Suppose that $\overline{\varphi}_{i}^{v}$ is isotropic for some $i\in\{1,\dots,r\}$. Then $\varphi_{i}$ is isotropic, by 3.10 $(b)$, and hence $\varphi$ is isotropic.
Suppose now that $\overline{\varphi}_{i}^{v}$ is anisotropic for $1\leqslant i\leqslant r$. Then $\varphi_{1},\dots,\varphi_{r}$ are anisotropic and represent only elements of ${K}^{\times}$ that lie in $v^{-1}(2vK)$. Consider $x\in K^{n}\smallsetminus\{0\}$ for $n=\mathsf{dim}(\varphi)$ and let $a=\varphi(x)$. As $\varphi_{1},\dots,\varphi_{r}$ are anisotropic and $\varphi\simeq c_{1}\varphi_{1}\perp\ldots\perp c_{r}\varphi_{r}$, we obtain that $a=c_{1}a_{1}+\ldots+c_{r}a_{r}$ for some $(a_{1},\dots,a_{r})\in K^{r}\smallsetminus\{0\}$ with $v(a_{1}),\dots,v(a_{r})\in 2vK\cup\{\infty\}$. For $1\leqslant i<j\leqslant r$, we have $v(c_{i})\not\equiv v(c_{j})\mod 2vK$, so either $v(c_{i}a_{i})\neq v(c_{j}a_{j})$ or $a_{i}=a_{j}=0$. As $(a_{1},\dots,a_{r})\neq 0$, we conclude that $v(a)=\mathsf{min}\{v(a_{i}c_{i})\mid 1\leqslant i\leqslant r\}<\infty$, whereby $a\neq 0$. This argument shows that $\varphi$ is anisotropic.
Having thus shown the equivalence for anisotropy, we will use it now to establish the equivalence for hyperbolicity. For $i\in\{1,\dots,r\}$, let $\psi_{i}$ denote the anisotropic part of $\varphi_{i}$ and let $m_{i}\in\mathbb{N}$ be such that $\varphi_{i}\simeq m_{i}\times\mathbb{H}\perp\psi_{i}$. Set $m=m_{1}+\ldots+m_{r}$ and $\psi=c_{1}\psi_{1}\perp\ldots\perp c_{r}\psi_{r}$. Then $\varphi\simeq m\times\mathbb{H}\perp\psi$. By 3.6, $\psi_{1},\dots,\psi_{r}$ are $v$-unimodular. It follows by the above that $\psi$ is anisotropic. Therefore $\varphi$ is hyperbolic if and only if $\psi$ is trivial, if and only if $\psi_{1},\dots,\psi_{r}$ are all trivial. Furthermore, for $1\leqslant i\leqslant r$, we have by part $(c)$ of 3.10 that $\overline{\varphi}_{i}^{v}$ is hyperbolic if and only if $\varphi_{i}$ is hyperbolic, which is if and only if the anisotropic part $\psi_{i}$ is trivial. This establishes the equivalence for hyperbolicity.
Consider now $m\in\mathbb{N}$. Note that $m\times\varphi\simeq c_{1}(m\times\varphi_{1})\perp\ldots\perp c_{r}(m\times\varphi_{r})$ and that $m\times\varphi_{i}$ is $v$-unimodular for $1\leqslant i\leqslant r$. Having this for all positive integers $m$, the equivalence for being torsion follows from that for hyperbolicity. ∎
Remark 3.12.
When restricting attention to complete discrete valuation rings and their fraction fields, most (if not all) of the results in this section are well-known. To the best of our knowledge these first appeared explicitly in Durfee’s doctoral research [9] for fields with a complete rank $1$ valuation. Inspection reveals that Durfee’s proofs (and other versions found later) mostly work without changes for general valued fields $(K,v)$ with $1+\mathfrak{m}_{v}\subseteq K^{\times 2}$ and $2\in\mathcal{O}_{v}^{\times}$, thus in particular for non-dyadic henselian valued fields, which would be sufficient for the purposes in the rest of this article. For our own and the reader’s convenience, we have included here an elementary and self-contained exposition which postpones adding the assumption $1+\mathfrak{m}_{v}\subseteq K^{\times 2}$ until the very end.
4. The generalised $u$-invariant
Let $K$ be a field with $\mathsf{char}(K)\neq 2$. Recall that we defined a quadratic form $\varphi$ over the field $K$ to be a torsion from if $m\times\varphi=\varphi\perp\ldots\perp\varphi$ is hyperbolic for some $m\in\mathbb{N}^{+}$. The general $u$-invariant of $K$ as defined by Elman and Lam in [11] is given by
$u(K)\,=\,\mathsf{sup}\{\mathsf{dim}(\varphi)\mid\varphi\mbox{ anisotropic torsion form over }K\}\,\in\mathbb{N}\cup\{\infty\}.$
Note that, if $K$ is nonreal, then every quadratic form over $K$ is a torsion form, and hence in this case the definition can be simplified to
$u(K)\,=\,\mathsf{sup}\{\mathsf{dim}(\varphi)\mid\varphi\mbox{ anisotropic formover }K\},$
which corresponds to how the $u$-invariant was originally defined by Lenz.
Let $v$ in the sequel be a valuation on $K$ with $\mathsf{char}(Kv)\neq 2$.
Proposition 4.1.
We have $u(K)\geqslant[vK:2vK]\cdot u(Kv)$, except possibly when $K$ is real, $Kv$ is nonreal and $u(Kv)$ is odd. In any case, $u(K)\geqslant[vK:2vK]\cdot 2\lfloor\mbox{$\frac{u(Kv)}{2}$}\rfloor$.
Proof.
See [5, Theorem 5.2]. ∎
The following statement and its argument are essentially known; see e.g. [27, Theorem 4].
Proposition 4.2.
Assume that $1+\mathfrak{m}_{v}\subseteq{K}^{\times 2}$. Then
$u(K)=[vK:2vK]\cdot u(Kv)\,.$
Proof.
In view of 3.10 $(a)$, the hypothesis implies that $K$ is real if and only if $Kv$ is real. In particular, 4.1 yields that
$u(K)\geqslant[vK:2vK]\cdot u(Kv)\,.$
It remains to show the converse inequality.
For this, we set $r=[vK:2vK]$ and $u=u(Kv)$ and assume without loss of generality that $r<\infty$. Note that $v^{-1}(2K)={\mathcal{O}}^{\times}_{v}{K}^{\times 2}$. We fix representatives $c_{1},\dots,c_{r}\in{K}^{\times}$ of the different classes of ${K}^{\times}/{\mathcal{O}}^{\times}_{v}{K}^{\times 2}$. Let $\varphi$ now be an anisotropic torsion form over $K$. We write $\varphi\simeq c_{1}\varphi_{1}\perp\ldots\perp c_{r}\varphi_{r}$ where $\varphi_{1},\dots,\varphi_{r}$ are $v$-unimodular forms over $K$. Then $\overline{\varphi}_{1}^{v},\dots,\overline{\varphi}_{r}^{v}$ are anisotropic torsion forms over $K$, by 3.11. Hence $\mathsf{dim}(\varphi_{i})=\mathsf{dim}(\overline{\varphi}_{i}^{v})\leqslant u$ for $1\leqslant i\leqslant r$, whereby $\mathsf{dim}(\varphi)=\sum_{i=1}^{r}\mathsf{dim}(\varphi_{i})\leqslant r\cdot u$. This shows that $u(K)\leqslant r\cdot u$. ∎
5. The generalised strong $u$-invariant
We are now prepared to study the $u$-invariant of a function fields in one variable. Let $K$ be a field with $\mathsf{char}(K)\neq 2$. As in [4, Sect. 5], we define
$\widehat{u}(K)\,=\,\mbox{$\frac{1}{2}$}\mathsf{sup}\,\{u(F)\mid F/K\mbox{ function field in one variable}\}\,\in\mbox{$\frac{1}{2}$}\mathbb{N}\cup\{\infty\}.$
We call this the generalised strong $u$-invariant of $K$. For nonreal fields, $\widehat{u}$ coincides with the ‘strong $u$-invariant’ defined in [15].
Proposition 5.1.
We have $u(K(X))\geqslant 2\,\mathsf{sup}\{u(L)\mid L/K\mbox{ finite field extension}\}$.
Proof.
See [5, Proposition 6.1]. ∎
Example 5.2.
In [7], an example is constructed of a field $K$ such that $u(L)=2$ holds for every finite field extension $L/K$, while there exists a pair of quadratic forms in $5$ variables over $K$ having no nontrivial common zero. As explained in [1, Introduction], this implies that $u(K(X))\geqslant 6$. Hence in this case we obtain that $\widehat{u}(K)\geqslant\frac{1}{2}u(K(X))\geqslant 3>2=\mathsf{sup}\{u(L)\mid L/K\mbox{ finite field extension}\}$.
The field $K$ is called quadratically closed if ${K}^{\times}={K}^{\times 2}$, and it is called hereditarily quadratically closed if ${L}^{\times}={L}^{\times 2}$ for every finite field extension $L/K$.
Proposition 5.3.
If $K(\sqrt{-1})$ is hereditarily quadratically closed, then $\widehat{u}(K)=1$, and otherwise $\widehat{u}(K)\geqslant 2$.
Proof.
If $K(\sqrt{-1})$ is hereditarily quadratically closed, then $u(F)=2$ for every function field in one variable $F/K$, by [12, Theorem], so in particular $\widehat{u}(K)=1$. Assume now $K(\sqrt{-1})$ is not hereditarily quadratically closed. Hence we may take a finite field extension $K^{\prime}/K(\sqrt{-1})$ such that $K^{\prime}$ is not quadratically closed. Then $u(K^{\prime})\geqslant 2$. As $K^{\prime}(X)$ is a function field in one variable over $K$ and $u(K^{\prime}(X)))\geqslant 2u(K^{\prime})\geqslant 4$, by 5.1, we conclude that $\widehat{u}(K)\geqslant 2$. ∎
Note that by its definition, $\widehat{u}(K)$ is always in $\frac{1}{2}\mathbb{N}\cup\{\infty\}$, but we do not know whether it is always in $\mathbb{N}$. We do not even have any example where $\widehat{u}(K)$ would be not a power of $2$.
Proposition 5.4.
Let $v$ be a valuation on $K$ with $\mathsf{char}(Kv)\neq 2$. Then
$\widehat{u}(K)\geqslant[vK:2vK]\cdot\widehat{u}(Kv)\,.$
Proof.
[…]
$[vK:2vK]\cdot u(E)=[wF:2wF]\cdot u(Fw)\leqslant u(F)\leqslant 2\widehat{u}(K)\,.$
Having this for all function fields in one variable $E/Kv$, the statement follows. ∎
The following proposition considers the behaviour of the invariant $\widehat{u}$ under algebraic field extensions. It recovers the fact that $\widehat{u}(K)$ is also an upper bound on $u(K^{\prime})$ for all finite field extensions $K^{\prime}/K$.
Proposition 5.5.
Let $K^{\prime}/K$ be an algebraic field extension. Then
$u(K^{\prime})\leqslant\widehat{u}(K^{\prime})\leqslant\widehat{u}(K)\,.$
Moreover, if $K^{\prime}/K$ is a direct limit of finite normal extensions of odd degree of $K$, then
$\widehat{u}(K^{\prime})=\widehat{u}(K)\,.$
Proof.
By 5.1, we have that $2u(K^{\prime})\leqslant u(K^{\prime}(X))\leqslant 2\widehat{u}(K^{\prime})$. Hence $u(K^{\prime})\leqslant\widehat{u}(K^{\prime})$.
Consider a function field in one variable $F^{\prime}/K^{\prime}$. Consider an anisotropic torsion form $\varphi$ over $F^{\prime}$. There exists a finitely generated subextension $F/K$ of $F^{\prime}/K$ such that $\varphi$ is defined over $F$ and is a torsion form over $F$. Moreover, we may choose $F$ to be transcendental over $K$. Then $F/K$ is a function field in one variable and $\varphi$ is an anisotropic torsion form over $F$. Hence $\mathsf{dim}(\varphi)\leqslant u(F)\leqslant 2\widehat{u}(K)$. Since this holds for any anisotropic torsion form $\varphi$ over $F^{\prime}$, it follows that $u(F^{\prime})\leqslant 2\widehat{u}(K)$. Having this for any function field in one variable $F^{\prime}/K^{\prime}$, we conclude that $\widehat{u}(K^{\prime})\leqslant\widehat{u}(K)$.
Assume now that $K^{\prime}/K$ is a direct limit of finite normal extensions of odd degree. It remains to show that in this case $\widehat{u}(K)\leqslant\widehat{u}(K^{\prime})$. Consider a function field in one variable $F/K$ and an anisotropic torsion form $\varphi$ over $F$. There exists a function field in one variable $F^{\prime}/K^{\prime}$ such that $F\subseteq F^{\prime}$ and $F^{\prime}=K^{\prime}F$. Since $K^{\prime}/K$ is a direct limit of finite normal extensions of odd degree, so is $F^{\prime}/F$. Therefore, Springer’s Theorem for odd degree extensions [17, Chap. VII, Theorem 2.7] yields that $\varphi$ remains anisotropic over $F^{\prime}$. Since it also remains a torsion form over $F^{\prime}$, we obtain that $\mathsf{dim}(\varphi)\leqslant u(F^{\prime})\leqslant 2\widehat{u}(K^{\prime})$. Since this holds for all anisotropic torsion forms $\varphi$ over $F$, we see that $u(F)\leqslant 2\widehat{u}(K^{\prime})$. Having this for all function fields in one variable $F/K$, we conclude that $\widehat{u}(K)\leqslant\widehat{u}(K^{\prime})$. ∎
Note in the context of 5.5 that, for a general algebraic extension $K^{\prime}/K$, one may have that $\widehat{u}(K^{\prime})<\widehat{u}(K)$: in view of 5.3, it suffices to have that $K(\sqrt{-1})$ is not hereditarily quadratically closed and to take for $K^{\prime}$ an algebraic closure of $K$.
Question 5.6.
Does $\widehat{u}(K^{\prime})=\widehat{u}(K)$ hold for every finite field extension $K^{\prime}/K$?
In most cases where we have any knowledge on the $u$-invariant of function fields over $K$, we have in particular that $u(K(X,Y))=4u(K)$.
Corollary 5.7.
Assume that $u(K(X,Y))=4u(K)$. Then $\widehat{u}(K)=u(K)$.
Proof.
By 5.5, we have $u(K)\leqslant\widehat{u}(K)$. To show the converse inequality, we need to show that $u(F)\leqslant 2u(K)$ for an arbitrary function field in one variable $F/K$. Given such $F/K$, there exist $x,y\in F$ such that $F/K(x,y)$ is a finite purely inseparable field extension. In particular, $K(x,y)$ contains a subfield $F_{0}$ which is isomorphic to $F$ and such that $F/F_{0}$ is a direct limit of odd degree extensions, and we obtain by using Springer’s Theorem for odd degree extensions [17, Chap. VII, Theorem 2.7] that $u(F)=u(F^{\prime})\leqslant u(K(x,y))\leqslant u(F)$, whereby $u(F)=u(K(x,y))$. We may thus assume that $F=K(x,y)$ for some $x,y\in F$. Then $F$ is the residue field of a discrete valuation on $K(X,Y)$. It follows by 4.1 and the hypothesis that $4u(K)=u(K(X,Y))\geqslant 2u(F)$, whereby $u(F)\leqslant 2u(K)$, as desired. ∎
By 5.1, we have $u(K(X))\geqslant 2u(K)$. This has an analogue for the strong $u$-invariant.
Proposition 5.8.
We have $\widehat{u}(K(X))\geqslant 2\widehat{u}(K)$.
Proof.
Consider a function field in one variable $F/K$. Then $F(X)/K(X)$ is function field in one variable and $u(F(X))\geqslant 2u(F)$, by 5.1. Therefore we have $2\widehat{u}(K(X))\geqslant u(F(X))\geqslant 2u(F)$. Having this for every function field in one variable $F/K$ yields the claimed inequality. ∎
Question 5.9.
Is $\widehat{u}(K(X))=2\widehat{u}(K)$?
For the $u$-invariant, the corresponding question would have a negative answer.
Example 5.10.
Let $K$ denote the quadratic closure of $\mathbb{Q}$. Choose an arbitrary proper finite field extension $L/K$ (e.g. $L=K(\sqrt[5]{6})$). Then $L$ is nonreal, and by [17, Chap. VII, Cor. 7.11] we have $|{L}^{\times}/{L}^{\times 2}|=\infty$. In particular $u(L)\geqslant 2$. It follows by 5.1 that $u(K(X))\geqslant 2u(L)\geqslant 4>2=2u(K)$.
One may wonder whether $\widehat{u}(K)$ is characterised in terms of the $u$-invariant by a single field extension of $K$. For example, is $u(F)\leqslant u(K(X))$ for every function field in one variable $F/K$? In other terms:
Question 5.11.
Is $\widehat{u}(K)=\frac{1}{2}u(K(X))$?
6. Function fields in one variable over a henselian valued field
We will now prove our main result and subsequently present various situations where it can be applied. We need a last step of preparation. For a rank-$1$ valuation $w$ on a field $F$, we denote by $F^{w}$ the corresponding completion. Let $K$ be a field with $\mathsf{char}(K)\neq 2$.
Lemma 6.1.
Let $v$ be a henselian rank-$1$ valuation on $K$ and $\mathsf{char}(Kv)\neq 2$. Let ${F/K}$ be a function field in one variable and $w$ a rank-$1$ valuation on $F$ with $\mathcal{O}_{v}\subseteq\mathcal{O}_{w}$. Then
$u(F^{w})=[wF:2wF]\cdot u(Fw)\leqslant[vK:2vK]\cdot 2\widehat{u}(Kv)\,.$
Proof.
Since $2\in{\mathcal{O}}^{\times}_{v}\subseteq{\mathcal{O}}^{\times}_{w}$, we have that $\mathsf{char}(Fw)\neq 2$. Note that $w$ extends naturally to a complete valuation $w^{\prime}$ on $F^{w}$ with $w^{\prime}F^{w}=wF$ and $F^{w}w^{\prime}=Fw$. Hence $w^{\prime}$ is a complete rank-$1$ valuation on $F$ and $\mathsf{char}(F^{w}w^{\prime})\neq 2$. It follows by [13, Theorem 1.3.1 (Hensel’s Lemma)] that $1+\mathfrak{m}_{w^{\prime}}\subseteq{(F^{w})}^{\times 2}$. The equality on the left follows now from 4.2 applied to $(F^{w},w^{\prime})$. It remains to show that $[wF:2wF]\cdot u(Fw)\leqslant[vK:2vK]\cdot 2\widehat{u}(Kv)$.
The situation where $w|_{K}$ is trivial will be treated below as case $(4)$. We first consider the situation where $w|_{K}$ is nontrivial. Then $\mathcal{O}_{v}\subseteq\mathcal{O}_{w}\cap K\subsetneq K$. Since $v$ has rank $1$, this implies that $w|_{K}$ is equivalent to $v$. We may then assume that $v=w|_{K}$. We follow the case distinction given in 2.2.
$(1)$ Suppose that $Fw/Kv$ is a finite field extension and $wF/vK$ is a finitely generated $\mathbb{Z}$-module with $\mathsf{rrk}(Fw/Kv)=1$. It follows by 2.1 that $[wF:2wF]=2\cdot[vK:2vK]$. By 5.5 we have that $u(Fw)\leqslant\widehat{u}(Kv)$. From this the claimed inequality follows.
[…]
$(3)$ Suppose that $Fw/Kv$ is an algebraic field extension and $wF/vK$ is a torsion group. Then $u(Fw)\leqslant\widehat{u}(Kv)$ and $[wF:2wF]\leqslant[vK:2vK]$ by 2.1. In this case we obtain even that $[wF:2wF]\cdot u(Fw)\leqslant[vK:2vK]\cdot\widehat{u}(Kv)$.
$(4)$ Assume finally that $w|_{K}$ is trivial. In this case $w$ is a discrete valuation, so we may assume that $wF=\mathbb{Z}$, and the residue field $Fw$ is a finite extension of $K$. Hence $v$ extends to a valuation $v^{\prime}$ on $Fw$, and as $v$ is henselian, so is $v^{\prime}$. It follows by 4.2 that $u(F^{w})=[wF:2wF]\cdot u(Fw)=2u(Fw)$ and $u(Fw)=[v^{\prime}(Fw):2v^{\prime}(Fw)]\cdot u((Fw)v^{\prime})$. Since the extension $(Fw)v^{\prime}/Kv$ is finite, we have that $u((Fw)v^{\prime})\leqslant\widehat{u}(Kv)$, by 5.5. Furthermore, $[v^{\prime}(Fw):vK]\leqslant[Fw:K]<\infty$, whence $[v^{\prime}(Fw):2v^{\prime}(Fw)]=[vK:2vK]$ by 2.1. Thus $u(Fw)=[vK:2vK]\cdot u((Fw)v^{\prime})\leqslant[vK:2vK]\cdot\widehat{u}(Kv)$. ∎
We come to our generalisation of [28, Theorem 3] and [4, Cor. 6.4]. The proof crucially depends on the local-global principle for isotropy from [23, Cor. 3.9], or more precisely, on the slightly extended version provided in [3, Theorem 4.4], where the valuation on the base field is assumed of rank $1$ and henselian, but not necessarily complete.
Theorem 6.2.
Let $v$ be a henselian valuation on $K$ with $\mathsf{char}(Kv)\neq 2$. Then
$\widehat{u}(K)=[vK:2vK]\cdot\widehat{u}(Kv).$
Proof.
By 5.4, we have $\widehat{u}(K)\geqslant[vK:2vK]\cdot\widehat{u}(Kv)$, so it remains to show the converse inequality.
We first consider the case where $v$ has rank $1$. Consider an arbitrary function field in one variable $F/K$ and an anisotropic torsion form $\varphi$ over $F$. We claim that $\mathsf{dim}(\varphi)\leqslant[vK:2vK]\cdot 2\widehat{u}(Kv)$. If $\mathsf{dim}(\varphi)\leqslant 2$, then this is trivial, because $\widehat{u}(Kv)\geqslant 1$. Suppose now that $\mathsf{dim}(\varphi)\geqslant 3$. Since $\varphi$ is anisotropic over $F$ and $v$ is henselian of rank $1$, it follows by [3, Theorem 4.4] that there exists a rank-$1$ valuation $w$ on $F$ with $\mathcal{O}_{v}\subseteq\mathcal{O}_{w}$ and such that $\varphi$ remains anisotropic over $F^{w}$. We obtain by 6.1 that $\mathsf{dim}(\varphi)\leqslant[vK:2vK]\cdot 2\widehat{u}(Kv)$, as claimed. Having this established for all function fields $F/K$ and all anisotropic torsion forms $\varphi$ over $F$, we conclude that $\widehat{u}(K)=[vK:2vK]\cdot\widehat{u}(Kv)$. We have now established that $\widehat{u}(K)=[vK:2vK]\cdot\widehat{u}(Kv)$ whenever $v$ is of rank $1$.
[…]
We now consider the general case without any assumption on the value group $vK$. Let $\mathbb{F}$ denote the prime field of $K$. Consider an arbitrary function field in one variable $F/K$ and an anisotropic torsion form $\varphi$ over $F$. We fix a positive integer $m$ such that $m\times\varphi$ is hyperbolic. We choose a finite subset $S$ of $F$ such that $\varphi$ is defined over $K(S)$ and is a torsion form over $K(S)$, and further such that $F=K(S)$ and $\mathbb{F}(S)/\mathbb{F}(S\cap K)$ is of transcendence degree $1$. For any subfield $F_{1}$ of $F$ containing $S$, we have $u(F_{1})\geqslant\mathsf{dim}(\varphi)$ because $\varphi$ is an anisotropic torsion form over $F_{1}$.
We set $K_{0}=\mathbb{F}(S\cap K)$. Since the absolute transcendence degree of $K_{0}$ is bounded by $|S|$, it follows by [13, Theorem 3.4.3] that $\mathsf{rk}(v|_{K_{0}})\leqslant|S|+1<\infty$. By 2.3, there exists a subfield $K_{1}$ of $K$ containing $K_{0}$ such that $vK_{1}/vK_{0}$ is torsion, $vK/vK_{1}$ is torsion-free, $Kv/K_{1}v$ is algebraic and purely inseparable and $v|_{K_{1}}$ is henselian. As the extension $Kv/K_{1}v$ is purely inseparable and ${\mathsf{char}(Kv)\neq 2}$, it is in particular a direct limit of normal extensions of odd degree. We conclude by 5.5 that $\widehat{u}(Kv)\leqslant\widehat{u}(K_{1}v)$. As the quotient group $vK/vK_{1}$ is torsion-free, we have $[vK_{1}:2vK_{1}]\leqslant[vK:2vK]$. Since $vK_{1}/vK_{0}$ is torsion, we conclude that $\mathsf{rk}(v|_{K_{1}})=\mathsf{rk}(v|_{K_{0}})<\infty$. Therefore, in view of the previous part, we have $\widehat{u}(K_{1})=[vK_{1}:2vK_{1}]\cdot\widehat{u}(K_{1}v)\leqslant[vK:2vK]\cdot\widehat{u}(Kv)$. Letting $F_{1}=K_{1}(S)$, we have that $F_{1}/K_{1}$ is a function field in one variable and obtain that $\mathsf{dim}(\varphi)\leqslant u(F_{1})\leqslant 2\widehat{u}(K_{1})\leqslant[vK:2vK]\cdot 2\widehat{u}(Kv)$. Having this for any anisotropic torsion form $\varphi$ defined over any function field in one variable $F/K$, we conclude that $\widehat{u}(K)\leqslant[vK:2vK]\cdot\widehat{u}(Kv)$. Since the other inequality also holds, we have established the equality in all generality. ∎
Remark 6.3.
Assume that $v$ is a complete rank-$1$ valuation on $K$ such that $\mathsf{char}(Kv)\neq 2$ and let $r=\mathsf{rrk}(v)$. In this setting, in [21, Theorem 4.1] (independently from this work) the bound $\widehat{u}(K)\leqslant 2^{r}\widehat{u}(Kv)$ is shown, which strengthens [23, Cor. 6.2], where the same bound is given for the special case where $vK\simeq\mathbb{Z}^{r}$, while otherwise achieving the weaker bound $\widehat{u}(K)\leqslant 2^{r+1}\widehat{u}(Kv)$.
[…]
We now recall three well-known constructions of valued fields $(K,v)$ with given residue field $Kv$ and a given value group $vK$, which provides us with examples where we can apply 6.2.
Example 6.4.
[…]
Example 6.5.
Let $k$ be a field with $\mathsf{char}(k)\neq 2$ and $\Gamma$ an ordered abelian group. Let $K$ denote the field of Hahn series with respect to $\Gamma$ and with coefficients in $k$. (The elements of $K$ are series $\sum_{\gamma\in S}a_{\gamma}t^{\gamma}$ with well-ordered support $S\subseteq\Gamma$.) Then $K$ carries a natural valuation $v$ with value group $\Gamma$ and residue field $k$, and $(K,v)$ is henselian, by [22, Lemma 2.40]. By 6.2, we obtain that $\widehat{u}(K)=[\Gamma:2\Gamma]\cdot\widehat{u}(k)$.
Example 6.6.
Consider a valued field $(k,v)$ with $\mathsf{char}(kv)\neq 2$ and value group $vk$ contained in $\mathbb{R}$ as an ordered group. In particular $\mathsf{rk}(v)\leqslant 1$. Let $G$ denote the divisible closure of $vk$, which is still a subgroup of $\mathbb{R}$, and in fact a $\mathbb{Q}$-linear subspace of $\mathbb{R}$. We consider the quotient space $\mathbb{R}/G$. Let $n\in\mathbb{N}$ and choose $\gamma_{1},\dots,\gamma_{n}\in\mathbb{R}$ such that $\overline{\gamma}_{1},\dots,\overline{\gamma}_{n}$ are $\mathbb{Q}$-linearly independent in $\mathbb{R}/G$. Let $K=k(X_{1},\dots,X_{n})$. By [13, Theorem 2.2.1], $v$ extends to a valuation $w$ on $K$ such that $w(X_{i})=\gamma_{i}$ for $1\leqslant i\leqslant n$ and with $Kw=kv$ and $wK\simeq vk\times\mathbb{Z}^{n}$. In particular, we have $[wK:2wK]=2^{n}\cdot[vk:2vk]$.
Let $(K^{\prime},w^{\prime})$ be either the henselization or the completion of $(K,w)$. Then in particular $(K^{\prime},w^{\prime})$ is a henselian valued field with $w^{\prime}K^{\prime}=wK\simeq vk\times\mathbb{Z}^{n}$ and $K^{\prime}w^{\prime}=Kw=kv$. By 6.2, one obtains that $\widehat{u}(K^{\prime})=2^{n}[vk:2vk]\cdot\widehat{u}(kv)$.
We now recall how examples of valuations to which 6.2 is applicable are obtained naturally from a classical geometric setting.
Example 6.7.
Let $(k,v)$ be a discretely valued field, and $\mathcal{O}_{v}$ its valuation ring. Let us take a regular arithmetic surface $\mathcal{X}$ over $\mathcal{O}_{v}$, i.e. a regular irreducible flat model of a regular irreducible curve $X$ over $k$. Let $D$ be an irreducible divisor in $\mathcal{X}$, and $\eta_{D}$ its generic point. Then $\mathcal{O}_{\mathcal{X},\eta_{D}}$ is a regular local ring of dimension $1$, hence a discrete valuation ring. It induces a discrete valuation $v_{1}$ on the function field $K$ of $X$, with residue field $\kappa(D)$-the function field of $D$. One can then take a closed point $x\in D$, and similarly, it induces a discrete valuation $v_{2}$ on $\kappa(D)$ of residue field $\kappa(x)$. By composing the valuations $v_{1}$ and $v_{2}$, we obtain a rank-$2$ valuation $w$ on $K$ with value group $\mathbb{Z}\times\mathbb{Z}$ and residue field $\kappa(x)$. Let $F$ denote its henselization. Then $\widehat{u}(F)=4\widehat{u}(\kappa(x))$ and $\kappa(x)/kv$ is a finite extension of $kv$. This can for instance be applied to models of $\mathbb{P}_{\mathbb{Q}_{p}}^{1}$ over $\mathbb{Z}_{p}$ for any prime number ${p\neq 2}.$
Example 6.8.
Let $k$ be a field in which $-1$ is not a square. Let us fix coordinate functions ${X,Y}$ on $\mathbb{P}_{k}^{2}$. Let $D=\pi^{-1}((X^{2}+1))$, the divisor of $\mathbb{P}_{k}^{2}$ induced by the fiber over $(X^{2}+1)$ of the projection $\pi:\mathbb{P}_{k}^{2}\rightarrow\mathbb{P}_{k}^{1}$, $(X,Y)\mapsto X$. We note that $D$ can be identified to $\mathbb{P}_{k(i)}^{1}$, where $i=\sqrt{-1}$ and $k(i)$ is the residue field of $(X^{2}+1)$ in $\mathbb{P}_{k}^{1}$. Its generic point $\eta_{D}$ corresponds to the ideal $(X^{2}+1)$ in $k[X,Y],$ so the local ring $\mathcal{O}_{\mathbb{P}_{k}^{2},\eta_{D}}$ is $k[X,Y]_{(X^{2}+1)}$, and it is a discrete valuation ring. Let $v_{1}$ be the discrete valuation it induces on $k(X,Y)$. The corresponding residue field is $k(i)(Y)$.
Now let us take a closed point $y$ of $D$. It is a divisor of $D$, and a codimension-$2$ point of $\mathbb{P}_{k}^{2}$. Let us choose e.g. $y=(Y^{2}+i)$. The local ring $\mathcal{O}_{\mathbb{P}_{k}^{1},y}=k(i)[Y]_{(Y^{2}+i)}$ is a discrete valuation ring. Let $v_{2}$ denote the corresponding valuation on $k(i)(Y)$, the fraction field of $k(i)[Y]_{(Y^{2}+i)}$. Note that its residue field is $k(\sqrt{-i})$. By composing $v_{1}$ and $v_{2}$ we obtain a rank-$2$ valuation $w$ on $k(X,Y)$ with value group $\mathbb{Z}\times\mathbb{Z}$ and residue field $k(\sqrt{-i})$. Let $F$ denote the henselization of $k(X,Y)$ with respect to $w$. Then, $\widehat{u}(F)=4\widehat{u}(k(\sqrt{-i})).$
This same procedure can be successively applied to higher-dimensional normal varieties which have successive normal fibrations.
We conclude by pointing out that in many cases when 6.2 is applied with a particular residue field $Kv$, on can obtain the precise value of $u(F)$ of every individual function field in one variable $F/K$.
Corollary 6.9.
Let $v$ be a henselian valuation on $K$ with $\mathsf{char}(Kv)\neq 2$. Let $n\in\mathbb{N}$ be such that $\widehat{u}(Kv)=2^{n}$. Then $\widehat{u}(K)=2^{n}\cdot[vK:2vK]$. Furthermore, if $u(\ell)=2^{n}$ for every finite field extension $\ell/Kv(\sqrt{-1})$, then for every function field in one variable $F/K$ we have $u(F)=2^{n+1}\cdot[vK:2vK]$.
Proof.
The first part is just a restatement of 6.2. To prove the second part, assume now that $u(\ell)=2^{n}$ for every finite field extension $\ell/Kv(\sqrt{-1})$. Consider a function field in one variable $F/K$. We choose $x\in F$ transcendental over $K$ and a discrete valuation $w$ on $F$ trivial on $K$ and with $w(x^{2}+1)>0$. Then $K(\sqrt{-1})\subseteq Fw$. We set $K^{\prime}=Fw$. Then $K^{\prime}/K$ is a finite field extension, and we obtain by 4.1 that $u(F)\geqslant 2u(K^{\prime})$. We denote by $v^{\prime}$ the unique extension of $v$ to $K^{\prime}$. Then $(K^{\prime},v^{\prime})$ is henselian and, by [13, Cor. 3.2.3], we have $[v^{\prime}K^{\prime}:vK]\cdot[K^{\prime}v^{\prime}:Kv]\leqslant[K^{\prime}:K]<\infty$. We conclude by 2.1 and 4.2, respectively, that
$[v^{\prime}K^{\prime}:2v^{\prime}K^{\prime}]=[vK:2vK]$ and $u(K^{\prime})=[v^{\prime}K^{\prime}:2v^{\prime}K^{\prime}]\cdot u(K^{\prime}v^{\prime}).$
As $K(\sqrt{-1})\subseteq Fw=K^{\prime}$, we have $Kv(\sqrt{-1})\subseteq K^{\prime}v^{\prime}$. Since $K^{\prime}v^{\prime}/Kv(\sqrt{-1})$ is a finite field extension, our assumption applies and yields that $u(K^{\prime}v^{\prime})=2^{n}$. Hence
$2^{n+1}\cdot[vK:2vK]=2u(K^{\prime})\leqslant u(F)\leqslant 2\widehat{u}(K)=2^{n+1}\cdot[vK:2vK]\,.$
Therefore $u(F)=2^{n+1}\cdot[vK:2vK]$. ∎
In the following cases, 6.9 provides directly the exact value of $u(F)$ for every function field in one variable $F/K$.
Example 6.10.
Let $n\in\mathbb{N}$, $v$ a henselian valuation on $K$ with $[vK:2vK]=2^{n}$ and $F/K$ be a function field in one variable.
  1. $(a)$
    If $Kv$ is algebraically closed with $\mathsf{char}(Kv)\neq 2$, then $u(F)=2^{n+1}$.
  2. $(b)$
    If $Kv$ is real closed, then $u(F)=2^{n+1}$.
  3. $(c)$
    If $Kv$ is a finite field with odd cardinality, then $u(F)=2^{n+2}$.
  4. $(d)$
    If $Kv=\mathbb{Q}_{p}$ for a prime number $p$, then $u(F)=2^{n+3}$.
Acknowledgments
We would like to thank David Leep for inspiring discussions including various pointers to the literature, in particular concerning Section 3. Karim Johannes Becher acknowledges support by the Bijzonder Onderzoeksfonds (project BOF-DOCPRO4, Nr. 49084, ‘Constructive proofs for quadratic forms’). Nicolas Daans gratefully acknowledges support by Czech Science Foundation (GAČR) grant 21-00420M, Charles University PRIMUS Research Programme PRIMUS/24/SCI/010, and Research Foundation - Flanders (FWO), fellowship 1208225N.